Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.25 coming 3 years from today. The dividend should grow rapidly - at a rate of 30% per year - during Years 4 and 5, but after Year 5, growth should be a constant 9% per year. If the required return on Computech is 13%, what is the value of the stock today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Year 3 dividend = 1.25

Year 4 dividend = 1.25 (1 + 30%) = 1.625

Year 5 dividend = 1.625 (1 + 30%) = 2.1125

Year 6 dividend = 2.1125 (1 + 9%) = 2.302625

Value at year 5 = D6 / required rate - growth rate

Value at year 5 = 2.302625 / 0.13 - 0.09

Value at year 5 = 2.302625 / 0.04

Value at year 5 = 57.56563

Value of stock = 1.25 / (1 + 0.13)3 + 1.625 / (1 + 0.13)4 + 2.1125 / (1 + 0.13)5 + 57.56563 / (1 + 0.13)5

Value of stock = $34.25